OpenStudy (anonymous):

14. For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept. Show your work.

OpenStudy (anonymous):

@alexrodriguez @Bill_Cipher

OpenStudy (anonymous):

@alexandrea.lane

OpenStudy (sbuck98):

Notice that when x = 0 the value of y is -3/1 so this line "cuts" the y axis at y=-3.00000 y-intercept = -3/1 = -3.00000 Calculate the X-Intercept : When y = 0 the value of x is -3/2 Our line therefore "cuts" the x axis at x=-1.50000 x-intercept = -3/2 = -1.50000 Calculate the Slope : Slope is defined as the change in y divided by the change in x. We note that for x=0, the value of y is -3.000 and for x=2.000, the value of y is -7.000. So, for a change of 2.000 in x (The change in x is sometimes referred to as "RUN") we get a change of -7.000 - (-3.000) = -4.000 in y. (The change in y is sometimes referred to as "RISE" and the Slope is m = RISE / RUN) Slope = -4.000/2.000 = -2.000

OpenStudy (anonymous):

ok. how

OpenStudy (sbuck98):

Read what I typed and it shows you how(:

OpenStudy (neonumbrella5115):

So first, find the y intercept, which is where the function crosses the y axis. You can do this by setting x as 0 and solving the equation to find a value for y. What do you get when you do this?

OpenStudy (sbuck98):

Solve y+2x+3 = 0 "y=mx+b" is the formula of a straight line drawn on Cartesian coordinate system in which "y" is the vertical axis and "x" the horizontal axis. In this formula : y tells us how far up the line goes x tells us how far along m is the Slope or Gradient i.e. how steep the line is b is the Y-intercept i.e. where the line crosses the Y axis The X and Y intercepts and the Slope are called the line properties. We shall now graph the line y+2x+3 = 0 and calculate its properties

OpenStudy (sbuck98):

^^ That's the first part, idk how it got deleted.

OpenStudy (anonymous):

thats a different problem though isnt it?